When determining the energy of a reaction where 0.250 mol of a substance with a ΔHcomb = -2250.0 kJ/mol reacts, we perform this calculation: −2250.0 kJ 1mol = x kJ 0.250 mol OR 0.250 mol −2250.0 kJ 1mol Why can’t we simply say that the energy of the reaction is -2250.0 kJ? (1 marks) ( )
1
Expert's answer
2020-06-04T10:31:44-0400

You should calculate the enthalpy per 0.250 mol, so you have to calculate ΔHcomb × n = -2250 × 0.250 = -562.5 (J).